CountingDaq 220513

From Moller Wiki
Jump to: navigation, search

Back to Main Page >> Data Acquisition Meetings >> Counting mode DAQ meetings

previous meeting << >> following meeting

Meeting information

 Meeting time:  12:00pm Eastern
 Zoom connection information
 Join ZoomGov Meeting
 Meeting ID: 160 380 0887
 Passcode: 903612
 US Tool free phone: (833) 568 8864
 Find your local number:


  • Activity updates
  • Thoughts about getting ready for 90% review
    • Taking the VETROC out of design, or just de-emphasise?
    • HDMI cable length tests; decision on MPD position; radiation environment evaluation for MPDs
    • Follow-up with grounding working group?
    • Beam monitor requirements for counting mode and how to instrument
    • Electrical safety evaluation of components; Electrical Safety Committee member for review?
  • General discussion


Recording on Zoom: [1]

Participants: R. Michaels, B. Moffit, C. Ghosh, D. Armstrong, M. Van Valkenburgh, C. Zorn, P. King

  • Paul introduces Morgan, an OU undergrad, who will be working on the parity mock data generator
  • Activity updates
    • Bob and Paul met with John Musson and Chad Seaton about what should be done to move forward on the digital beamline readout.
    • David reports that they have run with the Livermore low-energy EM list, and they see low flux from xrays below 100keV at the GEM locations. Cross-checks are needed, but it looks like the singles rates from the xrays will be no higher than the Moller+ep rate; for the xrays, they will only hit in a single layer. He plans to write this up shortly.
    • Discussion about the VETROC
      • Paul hasn't followed up with Mark P. to see if he would want the more precise timing
      • We may want to have some TDC channels anyway, such as comparing the GEM clock with the event trigger
      • David suggests if we want to put some PMTs into discriminator+VETROC, the only place he would imaging doing so would be the trigger scintillators
      • Bryan asks if we would really benefit from using the higher resolution.
    • Chandan has been testing a ring-6 detector with cosmics, and now at Mainz. The plan at UMass is to build a set of 1/14 of the main detector ring and test it with cosmics. Eventually they'd need two FADCs, right now they have one Hanjie had been using in the Solid test stand.
      • He estimates that it would be likely several months before they'd be ready for the 1/14-ring test.
      • UMass would also need some more HV to run 1/14-ring. Chandan will discuss with KK and then follow up.
    • Bob hasn't done any work on the test stand recently. He has ported the HV control to the rasberry-pi.
    • Bob asks where slow controls is included in the project. Do we have people taking care of EPICS for the detectors?
  • For the 90% review, we're still in a holding space until we better understand the beamline instrumentation
    • Bob suggests that we need to have some further testing done to know if the plans for readout changes are possible before proceeding
    • Does the DAQ group need more specification from tracking group as regards the beamline instrumentation? David will look at the beamline requirements document on SharePoint and let us know if he'd suggest changes